Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Mantovani at Sun Feb 17 23:17:26 MET 2002) Success
Cloudy all day

Torun (Kaz Borkowski at Mon Feb 18 10:05:50 MET 2002) Minor failures
There were some errors reported by FS but it seems they were not malignant. Tsys measured normally, BB7 out of order.

Jodrell_Bank (Alastair Gunn at Mon Feb 18 11:09:18 MET 2002) Minor failures
Bad count readback in VC04, therefore Tsys corrupted. Voltage still looks okay on VC04 so data should be good. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 18 11:09:38 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Effelsberg (A. Kraus at Mon Feb 18 11:54:38 MET 2002) Minor failures
Computer trouble: We lost data between UT1520-1546.

Onsala (John Conway at Tue Feb 19 08:25:07 MET 2002) Failure
Did not observe. Failure of recorder capstan motor.

Westerbork (Tony Foley at Tue Feb 19 14:39:17 MET 2002) Minor failures
scans 21:24 to 21:33 out of hour angle range. One telescope out of array (RTC was giving phase jumps)